Mass of flask and foil (g) = 63.4842 Mass of flask, foil, and condesned vapor (g) = 63.6995 Temperature of boiling water bath (C) = 95 degrees C Volume of flask (mL) = 134 Barometric pressre (mmHg) = 763.8 Room temperature (C) = 21.0 degrees C Vapor pressure at room temperature (mmHg) = 100 Find the following: 1) Uncorrected mass of condensed vapor (g) = ? 2). Corrected density of air (g/L) = ? 3). Average volume of flask (L) = ? 4). Initial mass of air in flask (g) = ? 5) Final partial pressure of air in flask (mmHg) = ? 6). Final mass of air in flask (g) = ? 7). Mass of air lost (g) = ? 8). Corrected mass of condensed vapor (g) = ? 9) Molecular weight (g/mol) = ?
Answer #1

Experiment -using ideal gas law to calculate molecular weight of vapour

1) Uncorrected mass of condensed vapor (g) = Mass of flask, foil, and condesned vapor (g) -Mass of flask and foil (g) =63.6995-63.4842=0.215g

2)Corrected density of air (g/L) =mass/Volume of flask (mL) =0.215g/0.134 L=1.6 g/L

3)Average volume of flask (L) = 134 ml*10^-3 L/ml=0.134ml

4)Initial mass of air in flask (g) = 0.215g

5)Final partial pressure of air in flask (mmHg) = 763.8 mmHg-100 mmhg=663.8 mmhg

6)Final mass of air in flask (g) calculation

no of moles of gas (vapour )escaped =PV/RT

P=763.8 mmHg*1atm/760mmhg=1 .005 atm

V=0.134L

T=95+273=368 C

n=1.005 atm*0.134L/0.0821 L atm/K.mol *368K=0.00446 moles

no of moles of air in the volume=n0=poV/RTo

po=100mmhg=100/760=0.131 atm

T=21+273=294 K

no=0.131 atm*0.134 L/0.0821L atm/K.mol *294K=7.27 *10^-4 moles

moles of liquid vapour=0.00446 moles-7.27 *10^-4 moles=0.00373 moles

7)Mass of air lost (g) = moles of air lost*molar mass of air=7.27 *10^-4 moles*28.964 g/mol=0.021 g

8)Corrected mass of condensed vapor (g) = 0.215g-0.021g=0.194 g

9) Molecular weight (g/mol) = mass of vapour /moles of vapour=0.194 g/0.00373 moles=52.01 g/mol